# Power and Play: Investigating “License to Critique” in Teams’ AI Ethics Discussions

### v1

@@ Abstract.
Past work has sought to design AI ethics interventions–such as checklists or toolkits–to help practitioners design more ethical AI systems. However, other work demonstrates how these interventions and the principles they’re based on may serve to instead limit critique to those addressed within the intervention, while rendering broader concerns illegitimate. In this paper, drawing on work examining how standards enact discursive closure and how power relations affect whether and how people raise critique, we recruit three corporate teams, and one activist team, each with prior context working with one another, to play a game designed to trigger broad discussion around AI ethics. We use this as a point of contrast to trigger reflection on their teams’ past discussions, examining factors which may affect their “license to critique” in AI ethics discussions. We then report on how particular affordances of this game may influence discussion, and find that the hypothetical context created in the game is unlikely to be a viable mechanism for real world change. We discuss how power dynamics within a group and notions of “scope” affect whether people may be willing to raise critique in AI ethics discussions, and discuss our finding that games are unlikely to enable direct changes to products or practice, but may be more likely to allow members to find critically-aligned allies for future collective action.
@@ 1. Introduction and Related Work
Many companies have Responsible AI guidelines, which often revolve around principles like Fairness, Accountability, and Transparency (Jobin et al., 2019). However, research finds that structural and systemic limits encoded in Silicon Valley logics, such as technosolutionism and the normalization of failure, limit those who seek to enact change (Metcalf et al., 2019).
@@
High-profile incidents illustrate these limits. In separate incidents, Meredith Whittaker (Conger and Wakabayashi, 2019) and Timnit Gebru (Metz and Wakabayashi, 2020) departed Google after seeking to organize against building military drone tech and the company’s handling of sexual harassment, and bias and environmental impacts of ever-larger language models, respectively. In a more recent case, so-called “godfather of AI” Geoffrey Hinton left Google in the spring of 2023, “so he can freely speak out about the risks of AI” (Metz, 2023) and not impact his employer while doing so. While different in numerous ways, these cases demonstrate that there are limits to the kind of direct ethical critique acceptable within tech companies.
@@
The prolific adoption of Responsible AI standards in companies (Jobin et al., 2019) may initially seem to legitimize workers as they raise ethical concerns (Madaio et al., 2020a). However, other work shows that they may, in fact, have the side effect of setting specific boundaries on what is and is not legitimate to raise as an ethical concern. In their work examining standards encouraging environmental sustainability in companies, Christensen et al. show how standards can enact “discursive closure,” which they understand in light of Deetz (Deetz, 1992) to mean legitimizing certain narrow kinds of employee critique while tacitly ruling others out of scope (Christensen et al., 2017). They suggest that a “license to critique” must be deliberately created to work against this limiting of discourse so that such standards can be flexible and enable discussion of concerns they do not specifically enumerate.
@@
Analogous arguments exist for the effect of Responsible AI standards. For example, Keyes et al. satirically argue that a system can be Fair, Accountable, and Transparent yet still “turn the elderly into high-nutrient slurry,” showing how certain harms may be outside of the scope of discursively closed principles (Keyes et al., 2019). Greene et al. show that the language of Responsible AI standards focuses scrutiny on the design of an AI system, but this focus thereby “reject[s] critiques of business practice” (Greene et al., 2019). In this sense, principles such as Fair, Accountable or Transparent that define responsible AI standards, or their focus on AI design rather than underlying business logics may set the discursive limits of the de facto language of AI ethics, thereby rendering concerns outside of this language less legitimate or intelligible, [^1] Indeed, this constrained language of AI ethics may operate similarly to the de facto language of “doing diversity”, in a way that similarly “conceals inequalities and neutralises histories of antagonism and struggle.” (Ahmed and Swan, 2006) both for workers in companies and for activists and policymakers seeking to influence companies’ actions. Put simply, if an ethical issue is not addressed by a guideline or standard, then it may be seen as out of bounds for critique, thereby limiting which issues can be discussed. In particular, standards and guidelines often seek to help practitioners build better AI systems (Greene et al., 2019), instead of helping them to critique the (business) ends to which they may be put, limiting critique to technical implementation instead of questioning if the system should be built at all.
@@ 2.1. Procedure
Given that we seek to examine how games affect AI ethics discussions rather than build a game ourselves, we choose to conduct the present study using Martelaro et al.’s What Could Go Wrong? a game in which groups of 4–5 participants discuss a series of AI applications and potential harms. Other games for AI ethics discussion exist (Ballard et al., 2019), but we choose this one because its source materials are readily available, can be easily adapted by adding cards, includes an online version for remote participants, [^2] https://github.com/nikmart/what-could-go-wrong-ai and because it is modeled on popular party games Apples to Apples or Cards Against Humanity, which may make it more easily understood by participants.
@@ 5. Discussion
Our results show that in game and in ordinary work, discussants seek to understand and display sensitivity to both the differentiated power and critical orientation of their discussion partners, which they use to modulate AI ethics issues they choose to raise and how they present them. When one’s boss or colleagues of unknown critical orientation are in the room, people may be less willing to raise critique. This echoes the work of James Scott, demonstrating how people employ “public transcripts” when those with power over them are present, but use more frank offstage talk when speaking to teammates they trust (Scott, 1990). Additionally, a variety of factors affect people’s perception of their own status, such as their seniority in the team or their proximity to engineering knowledge, in turn also affecting their willingness to raise critique.
@@ 6. Conclusion
In this paper, we examined how teams discuss AI ethics issues by drawing on past work examining how standards in other contexts enact discursive closure (Christensen et al., 2017), and on how power relations affect whether and how critique is raised (Scott, 1990). We recruit three corporate teams, and one activist team, each with prior context with one another, to play a game designed to trigger broad discussion around AI ethics, and firstly use this as a point of contrast to trigger reflection on their teams’ past discussions, examining organizational and team dynamics which may affect their “license to critique” in AI ethics discussion. We then report on how particular affordances of this game may influence discussion, paying particular attention to hypothetical games as a viable mechanism for real world change. We discuss how power dynamics in a group and notions of “scope” affect whether people may be willing to raise critique in AI ethics discussions, and our finding that games may not be able to lead to direct change, but may be more likely to allow members to find critically-aligned allies for future action.
@@ 7. Acknowledgments
We are grateful to our anonymous participants, Michael Madaio, Dawn Nafus, Lucy Suchman, and Julie Martinson Widder, who gave feedback on earlier drafts, as well as the support of the Digital Life Initiative at Cornell Tech.

### v2

@@ Abstract.
Past work has sought to design AI ethics interventions–such as checklists or toolkits–to help practitioners design more ethical AI systems. However, other work demonstrates how these interventions may instead serve to limit critique to that addressed within the intervention, while rendering broader concerns illegitimate. In this paper, drawing on work examining how standards enact discursive closure and how power relations affect whether and how people raise critique, we recruit three corporate teams, and one activist team, each with prior context working with one another, to play a game designed to trigger broad discussion around AI ethics. We use this as a point of contrast to trigger reflection on their teams’ past discussions, examining factors which may affect their “license to critique” in AI ethics discussions. We then report on how particular affordances of this game may influence discussion, and find that the hypothetical context created in the game is unlikely to be a viable mechanism for real world change. We discuss how power dynamics within a group and notions of “scope” affect whether people may be willing to raise critique in AI ethics discussions, and discuss our finding that games are unlikely to enable direct changes to products or practice, but may be more likely to allow members to find critically-aligned allies for future collective action.
@@ 1. Introduction and Related Work
Many companies have Responsible AI guidelines, which often revolve around principles like Fairness, Accountability, and Transparency (Jobin et al., 2019). However, research finds that structural and systemic limits encoded in Silicon Valley logics, such as technosolutionism and the normalization of failure, limit those who seek to enact change (Metcalf et al., 2019). High-profile incidents illustrate these limits. In separate incidents, Meredith Whittaker (Conger and Wakabayashi, 2019) and Timnit Gebru (Metz and Wakabayashi, 2020) departed Google after seeking to organize against building military drone tech and the company’s handling of sexual harassment, and bias and environmental impacts of ever-larger language models, respectively. Such cases demonstrate that there are limits to the kind of direct ethical critique acceptable within tech companies.
@@
The prolific adoption of Responsible AI standards in companies (Jobin et al., 2019) may initially seem to legitimize workers as they raise ethical concerns (Madaio et al., 2020a). However, other work shows that they may, in fact, have the side effect of setting specific boundaries on what is and is not legitimate to raise as an ethical concern. In their work examining environmental sustainability standards in companies, Christensen et al. show how such standards enact “discursive closure,” which they understand in light of Deetz (Deetz, 1992) to mean legitimizing certain narrow kinds of employee critique while tacitly ruling others out of scope (Christensen et al., 2017). They suggest that a “license to critique” must be deliberately created to work against this limiting of discourse so that such standards can be flexible and enable discussion of concerns they do not specifically enumerate.
@@
Analogous arguments exist regarding the effect of responsible AI standards. For example, Keyes et al. satirically argue that a system can be Fair, Accountable, and Transparent yet still “turn the elderly into high-nutrient slurry,” showing how certain harms may be outside of the scope of discursively closed principles (Keyes et al., 2019). Greene et al. show that the language of Responsible AI standards focuses scrutiny on the design of an AI system, but this focus thereby “reject[s] critiques of business practice” (Greene et al., 2019). In this sense, principles such as Fair, Accountable, or Transparent that define responsible AI standards, or their focus on AI design rather than underlying business logics may set the discursive limits of the de facto language of AI ethics, thereby rendering concerns outside of this language less legitimate or intelligible, [^1] Indeed, this constrained language of AI ethics may operate similarly to the de facto language of “doing diversity,” in a way that similarly “conceals inequalities and neutralises histories of antagonism and struggle.” (Ahmed and Swan, 2006) both for workers in companies and for activists and policymakers seeking to influence companies’ actions. Put simply, if an ethical issue is not addressed by a guideline or standard, then it may be seen as out of bounds for critique, thereby limiting which issues can be discussed. In particular, standards and guidelines often seek to help practitioners build better AI systems (Greene et al., 2019), instead of helping them to critique the (business) ends to which they may be put, limiting critique to technical implementation instead of questioning if the system should be built at all.
@@ 2.1. Procedure
Given that we seek to examine how games affect AI ethics discussions rather than build a game ourselves, we choose to conduct the present study using Martelaro & Ju’s What Could Go Wrong? a game in which groups of 4–5 participants discuss a series of AI applications and potential harms. Other games for AI ethics discussion exist (Ballard et al., 2019), but we choose this one because its source materials are readily available, can be easily adapted by adding cards, includes an online version for remote participants, [^2] https://github.com/nikmart/what-could-go-wrong-ai and because it is modeled on popular party games Apples to Apples or Cards Against Humanity, which may make it more easily understood by participants.
@@ 5. Discussion
Our results show that in the game and in ordinary work, discussants seek to understand and display sensitivity to both the differentiated power and critical orientation of their discussion partners, which they use to modulate AI ethics issues they choose to raise and how they present them. When one’s boss or colleagues of unknown critical orientation are in the room, people may be less willing to raise critique. This echoes the work of James Scott, demonstrating how people employ “public transcripts” when those with power over them are present, but use more frank offstage talk when speaking to teammates they trust (Scott, 1990). Additionally, a variety of factors affect people’s perception of their own status, such as their seniority in the team or their proximity to engineering knowledge, in turn also affecting their willingness to raise critique.
@@ 6. Conclusion
In this paper, we examined how teams discuss AI ethics issues by drawing on past work examining how standards in other contexts enact discursive closure (Christensen et al., 2017), and on how power relations affect whether and how critique is raised (Scott, 1990). We recruit three corporate teams, and one activist team, each with prior context with one another, to play a game designed to trigger broad discussion around AI ethics, and firstly use this as a point of contrast to trigger reflection on their teams’ past discussions, examining organizational and team dynamics which may affect their “license to critique” in AI ethics discussion. We then report on how particular affordances of this game may influence discussion, paying particular attention to hypothetical games as a viable mechanism for real-world change. We discuss how power dynamics in a group and notions of “scope” affect whether people may be willing to raise critique in AI ethics discussions. Our finding suggest that games may not be able to lead to direct change, but may be more likely to allow members to find critically-aligned allies for future action.
@@ 7. Acknowledgments
We are grateful to Wendy Ju, who helped develop the What Could Go Wrong? card game with Nikolas Martelaro. We thank Justin Deyo, Katie Ciez and Grace Klein at the Carnegie Mellon University Block Center for Technology and Policy for for hosting early pilot tests, and to student research assistants Sherry Chen, Eunbea Ji and Wan Tong Li for helping run these tests. We are grateful to our anonymous participants, Michael Madaio, Dawn Nafus, Lucy Suchman, and Julie Martinson Widder, who gave feedback on earlier drafts, as well as the support of the Digital Life Initiative at Cornell Tech.